MEASUREMENTS OF N_{2}- AND O_{2}-PRESSURE-INDUCED BROADENING AND PRESSURE-INDUCED SHIFTS FOR ^{16}O^{12}C^{32}S TRANSITIONS IN THE ν_{3} BAND

M.A. KOSHELEV, M.Yu. TRETYAKOV, Institute of Applied Physics, Russian Academy of Sciences, Nizhnii Novgorod, Russia 603950; \underline{R.M. LEES}, LI-HONG XU, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5.

Nitrogen and oxygen pressure-broadening and pressure-induced shift coefficients for 42 transitions of ^{16}O^{12}C^{32}S with quantum number m from -25 to 49 in the P and R branches of the ν_{3} band at 2062 cm^{−1} have been measured at room temperature using the UNB high-resolution tunable diode laser spectrometer. Significant dependence of the broadening coefficients on rotational quantum number is observed; the low-m limits are approximately 0.110 and 0.097 cm^{−1}/atm for N_{2} and O_{2} broadening, respectively, decreasing to 0.085 and 0.072 cm^{−1}/atm by m = 50. Air-broadening and shift parameters have also been calculated from the N_{2} and O_{2} measurements. The results show good consistency with previous measurements for the ν_{1} and 2ν_{3} bands but a slight deviation at high m from the coefficients for the ν_{3} band reported in the GEISA and HITRAN databases.